Shift report 20010927 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Sep 27 2001 - 00:03:56 EDT

  • Next message: Apache: "Shift report 20010927 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Sep 27 00:03:56 2001, Shift: 16:00-24:00
    Supervisor: JH Lee, with: 
    Log book#9 pages 51 - 54
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam/Machine:
       - Cogged Beam until 17:50 
       - New store at 23:00, but dumped at 23:30 due to 
         a cryogenic problem
    
    2. Data: 
       - Finished MRS 4A at 90deg (2M trigger1)  
         Starting MRS 4B at 90deg for 2M trigger1
       - FS 1/3A Field at 3deg: 2M trig1 so far 
         3M more to go
     
    3. Developments:
       - Vertex trigger  (trigger5) timing has been adjusted (FV)
       - Rate study for trigger 5 
         Trigger1:Trigger3:Trigger4:Trigger5=3:1:7:1
         (at ZDC=90/sec)
         triggger3 and trigger5 overlap ~30%
       - Run 4973 was taken only with Trigger5: 27350 Trigger5
    
    4. No new detector problem
       - T3,T4,T5 current (at ZDC~300/sec FS@3)
         T3: 31,26,34
         T4: 24,19,16
         T5: 41,22,23
       
         
              
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Sep 27 2001 - 00:04:29 EDT